OpenStudy (anonymous):

A right triangle has a hypotenuse of 155 centimeters and a leg of 78 centimeters. What is the length of the other leg?

OpenStudy (anonymous):

hypotenuse ^2=side1^2+side^2 u can proceed

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

is it 134? or 133?

OpenStudy (anonymous):

155^2=7^2+c^2 fing c=?...

OpenStudy (anonymous):

yes, is it 134 or 133?

OpenStudy (amistre64):

if thoses are the only options .. none of the above?

OpenStudy (amistre64):

lol, i did 7^2 not 78^2

OpenStudy (anonymous):

answer is 133.9 =134

OpenStudy (anonymous):

A right triangle has a hypotenuse of 50 feet and a leg of 14 feet. Which is the length of the other leg?